OpenStudy (anonymous):

A fisherman travels downstream at full speed to his favorite fishing spot. The stream is running at a rate of 5 miles per hour, and he takes an hour to get to the fishing hole. He catches 5 fish and calls it a day. He takes twice as long at full speed to get back to the dock going against the current. How fast will his fishing boat go?

OpenStudy (anonymous):

It took him 2 hours to get back, because it took him an hour to get there and 2ce the time to get back.

OpenStudy (anonymous):

how fast was the boat going though

OpenStudy (anonymous):

oops i missed the question he goes at 2.5 miles per hour

OpenStudy (anonymous):

because it has the same distance and 2 times the time it is going at 1/2 the speed. if he was going at 5 miles per hour down and 1/2 that going back he goes at 2.5 miles per hour

OpenStudy (anonymous):

haha thanks

OpenStudy (anonymous):

let its speed is x, the distance is L, so (x+5)*1=L (x-5)*2=L so(x+5)=(x-5)*2 x=15

OpenStudy (anonymous):

AndrewNJ your answer was right